OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (anonymous):

lack of something likes = or + or else. check the problem please

OpenStudy (anonymous):

that's it actually it says nothing else

OpenStudy (anonymous):

even between sin2x and 2x?

OpenStudy (anonymous):

so what does sin 2 2x mean?

OpenStudy (anonymous):

cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x there sorry! :)

OpenStudy (anonymous):

ok. edit the question in case there is someone stop by and he has a good idea.I ll try

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

break 2 = 1+1 . and you have 1- 2sin^2(2x) = cos4x and 1 -2sin^2 x = cos 2x . and you have it done. start from left side hope this helps

OpenStudy (anonymous):

sorry start from right side

OpenStudy (anonymous):

cos 4x + cos 2x = 2-2 sin^2 x - 2 sin^2 x cos 4x + cos 2x = 1-2 sin^2 x 1-2 sin^2 x cos 4x + cos 2x = cos 4x + cos 2x ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you! :)